### v1

@@ 1 Introduction
Surrogate models simulate aerodynamic characteristics without requiring extensive CFD simulations by utilizing data from high-fidelity simulations or experiments and applying various mathematical and statistical techniques [6]. These models facilitate rapid iteration and optimization during the design process by significantly reducing computational resources and time [7, 8]. A common approach to building aerodynamic surrogate models involves using low-dimensional reduction techniques followed by regression modeling. Dimensionality reduction methods are first applied to high-dimensional aerodynamic data to capture the most significant features with fewer dimensions. The reduced-dimensional data is then used to train regression models, establishing a relationship between the simplified input space and the aerodynamic outputs. Among the regression models explored in aerodynamic surrogates some options are K-Nearest Neighbors (kNN) [9], Gaussian Process Regression (GPR) [10], polynomial chaos-kriging [11], and architectures based on Neural Networks (NN) [12], among others.
@@
For dimensionality reduction, techniques like Proper Orthogonal Decomposition (POD) and Principal Component Analysis (PCA) are commonly used. POD is employed for reduced-order modeling of complex problems [13], within a Galerkin projection framework (POD-Galerkin) [14], combined with a CFD flux residual minimization scheme [15], or coupled with an interpolation method [16, 17]. The use of POD (or PCA) in surrogate modeling is extensively explored [18, 19, 20]; however, this linear technique has limitations in capturing non-linear complexities. Although POD has been applied to non-linear problems, its underlying assumption that the input data lie in a low-dimensional, linear subspace makes it insufficient for highly non-linear features. In aerodynamics, accurately predicting shockwaves in the transonic flow regime poses a severe challenge for POD-based methods because such flows result in sharp, discontinuous changes in flow variables like pressure, temperature, density, and velocity across the shockwave.
@@
To address the limitations of linear assumptions in POD, manifold learning techniques offer a viable alternative. Manifold learning, a branch of statistical learning, aims to recognize the topologically closed surface (the manifold) over which the data lie on or near it [21]. The manifold is a geometric representation of the intrinsic relationships that connect data samples. Examples include Isomap, Multi-dimensional Scaling (MDS), and Local Linear Embedding (LLE).Surprisingly, the application of manifold learning in aeronautics is not widely exploited. Franz et al. [22] combined, for the first time, Isomap with a thin plate spline regressor into a surrogate model to predict pressure distribution ($C_{p}$)on a wing in the transonic regime. Similarly, Castellanos et al. [23] combined Isomap with a NN, improving the surrogate’s performance in predicting shockwaves. Isomap has also been used for adaptive sampling [24], leading to a method that iteratively generates samples based on a low-dimensional manifold of the training data. Recently, LLE has been applied to develop a multi-fidelity surrogate model for efficient aerodynamic predictions [25]. The challenge of these techniques is that they require additional regression algorithms to decode the low-dimensional manifold back to the original dimension of the input data, adding further complexity to the surrogate [26]. Other alternatives to classic POD (or PCA) and manifold learning strategies include the Gauss-Newton with Approximated Tensors (GNAT), a projection-based framework related to residual minimization [27], or the Discrete Empirical Interpolation Method (DEIM) [28], which rely on a classical POD-Galerkin approach while allowing the evaluation of non-linear terms of the governing equations with an additional POD basis. However, their application to aerodynamic surrogate modeling still needs to be well-established.
@@
Recent advances in neural networks provide a promising framework for developing surrogate models that address the limitations of previous approaches [7, 29, 8, 30, 31, 32, 33]. Aerodynamic surrogates based on graph neural networks (GNNs) have gained interest due to their ability to efficiently handle unstructured data and focus on local features, as demonstrated by Hines and Bekemeyer [34] in predicting aircraft surface pressure distributions. Additionally, neural networks are also employed for dimensionality reduction. In particular, autoencoders (AEs) [35, 36, 37, 38] have been increasingly used in aerodynamic modeling, effectively addressing the limitations of POD and manifold learning by non-linear transformations of the data into compact low-dimensional spaces while providing a global encoding-decoding mechanism.
@@
Among the different AE architectures, Variational Autoencoders (VAEs) [39] have proven effective in encoding the spatial information of fluid flows into non-linear low-dimensional latent spaces [40, 41]. Unlike traditional AEs, VAEs incorporate a probabilistic approach for representing observations in the latent space by adding an extra loss term specific to the latent variables. This loss term allows VAEs to learn a more continuous latent space that better fits posterior regressor models. However, a notable limitation of these models is the lack of orthogonality, a feature inherent in classical linear decomposition techniques such as POD, which can result in several entangled features within a single latent variable, complicating the interpretation of the latent space. To address this limitation, the $\beta$-Variational Autoencoder ($\beta$-VAE) architecture [42] has been proposed. This architecture modifies the VAE’s loss function by incorporating a regularization parameter, $\beta$, which balances reconstruction accuracy with regularization and latent space disentanglement [43]. The $\beta$ parameter is carefully selected to ensure that the latent space representation is nearly orthogonal while minimizing any increase in reconstruction error. The potential of these architectures is outlined in various studies where different Reduced Order Models (ROMs) were developed using this architecture to describe the evolution of turbulent flow fields [44, 45, 46]. These studies report excellent performance in reconstructing turbulent flows and have found that the latent space features in these architectures represent the most energetic POD modes. These recent findings suggest that the $\beta$-VAE can capture non-linear modes representing the most energetic or representative flow features [45, 46]. Consequently, these architectures hold significant potential for developing physically interpretable surrogate models [36].
@@
Notable applications of AE-based architectures in surrogate modeling for aerodynamics include the recent use of graph convolutional autoencoders for steady-state aerodynamics of transonic aircraft [47, 48]. Specifically for $\beta$-VAE models, Kang et al. [36] present a study of the extraction of information-intensive latent variables (LVs) from two-dimensional airfoil flow data and explores how these low-dimensional LVs correlate with the physical properties of the flow. This work illustrates the ability of the $\beta$-VAE models to discard redundant LVs as a result of proper tuning of the $\beta$ hyperparameter, and the argument is made that this behavior is indicative of the physical awareness of the resulting model. In a related approach, Moni et al. [49] used DEIM to non-linearly derive a reduced set of variables from flows around airfoils and wings. The reduced set of variables was then used to train an autoencoder with clustering (AE-C), to obtain a disentangled low-dimensional representation of the data with a prescribed number of clusters. In the clustered space, a nearest-neighbor algorithm is used to select the nearest point, which is then used to reconstruct the flow field prediction.
@@
Due to the potential of these methods and the need to further analyze their performance in 3D flows, this study leverages the physically interpretable and information-rich latent variables (LVs) derived from $\beta$-VAEs for efficient surrogate modeling of aerodynamic pressure distribution. Using a comprehensive database of CFD simulations, we demonstrate the ability of $\beta$-VAEs to extract key aerodynamic features autonomously. Our proposed surrogate modeling framework integrates $\beta$-VAEs with PCA pre-processing to compress high-dimensional input data without information loss. This novel approach involves a two-step dimensionality reduction, leading to more efficient $\beta$-VAE training. Then, GPR maps the physical ROM input variables to the $\beta$-VAE latent space points. Finally, this work includes a final fine-tuning process that reduces the dependence on the $\beta$ hyperparameter, allowing for easier model tuning. This integration and the fine-tuning process improve the accuracy and generalization capability of the model, marking a significant advancement in aerodynamic surrogate modeling using autoencoders.
@@ 2 Database and test case: XRF1 wing
The selected test case comprises a database of CFD simulations of the XRF1 model. This model, provided by Airbus™, is a research test case for showcasing various technologies applied to long-range, wide-body aircraft. The database was developed within the Group for Aeronautical Research and Technology in Europe (GARTEUR) AD/AG60 research project. Further information on the XRF1 research model can be found in Pattinson and Herring [50].
@@
Aerodynamic data were derived from Reynolds Averaged Navier-Stokes (RANS) simulations of the complete aircraft model to ensure realistic conditions through the interaction of various aerodynamic subsystems. The dataset focuses on the dimensionless pressure coefficient ($C_{p}$) distribution over the wing for different flight conditions (see an example in Figure 1b). It comprises 435 distinct flight conditions computed using the inviscid DLR TAU solver [51] at a fixed Reynolds number ($Re=2.5\times 10^{7}$). The flight conditions span the entire flight envelope of the aircraft, with Mach numbers ($M$) ranging from 0.5 to 0.96 and angles of attack ($\alpha$) from $0^{\circ}$ to $11.5^{\circ}$ (see Figure 1a).
@@
Figure 1: (a) Flight envelope with the tested flight conditions split in training set ( ) and test set ( ). (b) Example of $C_{p}$ distribution over the wing for $\left(\alpha,M\right)=\left(9.25^{\circ},0.7\right)$. Three planes are highlighted at wingspan percentages $\eta=0.1,0.5,0.9$    ( ).
@@
The 435 flow solutions in the database are randomly split into two sets: a training set (70% of the database, 305 cases) used to train the surrogate model [( ) in Figure 1a], and a test set (30% of the database, 130 cases) to evaluate the model’s ability to predict $C_{p}$ within the entire design space [( ) in Figure 1a]. The cases in the test set are excluded from the training process. For visualization, three cases from the test set are chosen: low complexity $\left(\alpha_{1},M_{1}\right)=\left(2.0^{\circ},0.504\right)$, intermediate complexity $\left(\alpha_{2},M_{2}\right)=\left(5.75^{\circ},0.704\right)$, and high complexity $\left(\alpha_{3},M_{3}\right)=\left(9.75^{\circ},0.904\right)$, the latter featuring shockwaves on the upper wing surface.
@@
The fixed load factor $n_{z}$ ($=L/W$, where $L$ is lift and $W$ is weight) of 2.5g poses an additional challenge for the prediction model due to significant wing deformation under high loads. This condition represents the limit load factor for transport category aircraft per airworthiness regulations. The wing undergoes structural deformation at such conditions but returns to its normal shape once the load factor returns to $n_{z}=1$ (typical cruise condition). Aerodynamically, the deflection of the wing geometry prevents shockwave formation on the upper surface, though significant $C_{p}$ changes and separation regions occur within the flight envelope. This condition is critical for surrogate model development due to its relevance to certification requirements.
@@
Based on the load factor condition, we only focus on the upper wing surface since the lower surface does not challenge the predictive models, as previously seen in Castellanos et al. [23]. Additionally, discarding the lower surface avoids the numerical artifacts in the model due to irregularities in the mesh around the engine pylon. Consequently, each $C_{p}$ sample is defined on an unstructured mesh of $q=49574$ grid points. Each $C_{p}$ distribution is an observation (point) in the high-dimensional space $\mathbb{R}^{q}$, where each dimension (feature) corresponds to a $C_{p}$ point. Let $\mathbf{X}\in{\mathbb{R}}^{q\times n}$ represent the data matrix, with $\mathbf{x}_{i}\in\mathbb{R}^{q}$ as each row ($C_{p}$ for $i=1,\ldots,n$). The dataset in $\mathbf{X}$ is inherently complex, challenging the extraction of a reduced set of meaningful coordinates that capture the main flow characteristics. Hereafter, data samples are denoted as $\mathbf{x}$, the dataset as $\mathbf{X}$, autoencoder reconstructions as $\mathbf{\tilde{x}}$, and surrogate model predictions as $\mathbf{\hat{x}}$. For clarity in the figures, the pressure distribution coefficient and its reconstructions and predictions are denoted as $C_{p}$, $\tilde{C}_{p}$, and $\hat{C}_{p}$, respectively.
@@ 3.1 $\beta-$variational autoencoder
Autoencoders are a recent field of research focused on using neural networks to derive meaningful, low-dimensional representations from high-dimensional data, leveraging the non-linear nature of neural networks and their versatility in handling various types of data. They are an unsupervised learning strategy that aims to learn a compressed representation (latent space) of the training data by encoding it into a lower-dimensional space through an encoder network and, then reconstructing the input from this latent representation using a decoder network.
We denote as $\mathbf{z}_{i}$ the $i$-th component of $\mathbf{Z}\in\mathcal{Z}$, being $\mathcal{Z}\subset\mathbb{R}^{d}$ the latent space of dimensionality $d$. The encoder network maps the input data, $\mathbf{x}$ in the data space $\mathcal{X}\subset\mathbb{R}^{q\times n}$, to the latent space ($E:\mathcal{X}\rightarrow\mathcal{Z}$) and the decoder network reconstructs the input data from the latent representation ($D:\mathcal{Z}\rightarrow\mathcal{X}$). The data is compressed into the latent space in the bottleneck between the encoder and decoder networks. Let $\tilde{\mathbf{x}}=(D\circ E)(\mathbf{x})$ be the output of the autoencoder, which is trained to minimize a reconstruction loss function $\mathcal{L}_{rec}$, to drive the reconstructed output data to match the input data. The standard autoencoder loss is given by the mean squared error (MSE),
$\mathcal{L}_{rec}=\lVert\mathbf{x}-(D\circ E)(\mathbf{x})\rVert_{2}^{2},$ (1)
@@
where $\odot$ refers to the element-wise product.
@@
In this paper, using $\beta$-VAEs to map pressure distribution fields enables the model to leverage a non-linear dimensionality reduction method with increased fidelity over a classical linear method.

### v2

@@ 1 Introduction
Surrogate models simulate aerodynamic characteristics without requiring extensive CFD simulations by utilizing data from high-fidelity simulations or experiments and applying various mathematical and statistical techniques [6]. These models facilitate rapid iteration and optimization during the design process by significantly reducing computational resources and time [7, 8]. A common approach to building aerodynamic surrogate models involves using low-dimensional reduction techniques followed by regression modeling. Dimensionality reduction methods are first applied to high-dimensional aerodynamic data to capture the most significant features with fewer dimensions. The reduced-dimensional data is then used to train regression models, establishing a relationship between the simplified input space and the aerodynamic outputs. Among the regression models explored in aerodynamic surrogates, some options are K-Nearest Neighbors (kNN) [9], Gaussian Process Regression (GPR) [10], polynomial chaos kriging [11], and architectures based on Neural Networks (NN) [12], among others.
@@
For dimensionality reduction, techniques like Proper Orthogonal Decomposition (POD) and Principal Component Analysis (PCA) are commonly used. POD is used for reduced-order modeling of complex problems [13], within a Galerkin projection framework (POD-Galerkin) [14], combined with a CFD flux residual minimization scheme [15], or coupled with an interpolation method [16, 17]. The use of POD (or PCA) in surrogate modeling is extensively explored [18, 19, 20]; however, this linear technique has limitations in capturing non-linear complexities. Although POD has been applied to nonlinear problems, its underlying assumption that the input data lie in a low-dimensional, linear subspace makes it insufficient for highly nonlinear features. In aerodynamics, accurately predicting shockwaves in the transonic flow regime poses a severe challenge for POD-based methods because such flows result in sharp, discontinuous changes in flow variables like pressure, temperature, density, and velocity across the shockwave.
@@
To address the limitations of linear assumptions in POD, manifold learning techniques offer a viable alternative. Manifold learning, a branch of statistical learning, aims to recognize the topologically closed surface (the manifold) over which the data lie on or near it [21]. The manifold is a geometric representation of the intrinsic relationships that connect data samples. Examples include Isomap, Multi-dimensional Scaling (MDS), and Local Linear Embedding (LLE). Surprisingly, the application of manifold learning in aeronautics is not widely exploited. Franz et al. [22] combined, for the first time, Isomap with a thin plate spline regressor into a surrogate model to predict the pressure distribution ($C_{p}$) on a wing in the transonic regime. Similarly, Castellanos et al. [23] combined Isomap with a NN, improving the surrogate’s performance in predicting shockwaves. Isomap has also been used for adaptive sampling [24], leading to a method that iteratively generates samples based on a low-dimensional manifold of the training data. Recently, LLE has been applied to develop a multi-fidelity surrogate model for efficient aerodynamic predictions [25]. The challenge of these techniques is that they require additional regression algorithms to decode the low-dimensional manifold back to the original dimension of the input data, adding further complexity to the surrogate [26]. Other alternatives to classic POD (or PCA) and manifold learning strategies include Gauss-Newton with Approximated Tensors (GNAT), a projection-based framework related to residual minimization [27], or the Discrete Empirical Interpolation Method (DEIM) [28], which rely on a classical POD-Galerkin approach while allowing the evaluation of nonlinear terms of the governing equations with an additional POD basis. However, their application to aerodynamic surrogate modeling still needs to be well established.
@@
Recent advances in neural networks provide a promising framework for developing surrogate models that address the limitations of previous approaches [7, 29, 8, 30, 31, 32, 33]. Aerodynamic surrogates based on graph neural networks (GNNs) have gained interest due to their ability to efficiently handle unstructured data and focus on local features, as demonstrated by Hines and Bekemeyer [34] in predicting aircraft surface pressure distributions. In addition, neural networks are also employed for dimensionality reduction. In particular, autoencoders (AE) [35, 36, 37, 38] have been increasingly used in aerodynamic modeling, effectively addressing the limitations of POD and manifold learning by non-linear transformations of the data into compact low-dimensional spaces while providing a global encoding-decoding mechanism.
@@
Among the different AE architectures, Variational Autoencoders (VAE) [39] have been proven to be effective in encoding the spatial information of fluid flows into low-dimensional non-linear latent spaces [40, 41]. Unlike traditional AEs, VAEs incorporate a probabilistic approach for representing observations in the latent space by adding an extra loss term specific to the latent variables. This loss term allows VAEs to learn a more continuous latent space that better fits posterior regressor models. However, a notable limitation of these models is the lack of orthogonality, a feature inherent in classical linear decomposition techniques such as POD, which can result in several entangled features within a single latent variable, complicating the interpretation of the latent space. To address this limitation, the $\beta$-Variational Autoencoder ($\beta$-VAE) architecture [42] has been proposed. This architecture modifies the VAE’s loss function by incorporating a regularization parameter, $\beta$, which balances reconstruction accuracy with regularization and latent space disentanglement [43]. The $\beta$ parameter is carefully selected to ensure that the latent space representation is nearly orthogonal while minimizing any increase in reconstruction error. The potential of these architectures is outlined in various studies where different Reduced Order Models (ROMs) were developed using this architecture to describe the evolution of turbulent flow fields [44, 45, 46]. These studies report excellent performance in reconstructing turbulent flows and have found that the latent space features in these architectures represent the most energetic POD modes. These recent findings suggest that $\beta$VAE can capture nonlinear modes that represent the most energetic or representative flow characteristics [45, 46]. Consequently, these architectures hold significant potential for developing physically interpretable surrogate models [36].
@@
Notable applications of AE-based architectures in surrogate modeling for aerodynamics include the recent use of graph convolutional autoencoders for steady-state aerodynamics of transonic aircraft [47, 48], although they can be extended to other fields such as aeroacoustic airfoil shape optimization [49] or modeling three-dimensional unsteady flows [50]. Specifically for $\beta$-VAE models, Kang et al. [36] present a study of the extraction of information-intensive latent variables (LVs) from two-dimensional airfoil flow data and explores how these low-dimensional LVs correlate with the physical properties of the flow. This work illustrates the ability of the $\beta$-VAE models to discard redundant LVs as a result of proper tuning of the $\beta$ hyperparameter, and the argument is made that this behavior is indicative of the physical awareness of the resulting model. In a related approach, Moni et al. [51] used DEIM to non-linearly derive a reduced set of variables from flows around airfoils and wings. The reduced set of variables was then used to train an autoencoder with clustering (AE-C), to obtain a disentangled low-dimensional representation of the data with a prescribed number of clusters. In the clustered space, a nearest-neighbor algorithm is used to select the nearest point, which is then used to reconstruct the flow field prediction.
@@
Due to the potential of these methods and the need to further analyze their performance in 3D flows, this study leverages the physically interpretable and information-rich LVs derived from $\beta$-VAEs for efficient surrogate modeling of aerodynamic pressure distribution. Using a comprehensive database of CFD simulations, we demonstrate the ability of $\beta$-VAEs to extract key aerodynamic features autonomously. Our proposed surrogate modeling framework integrates $\beta$-VAEs with PCA preprocessing to compress high-dimensional input data without loss of information. This novel approach involves a two-step reduction in dimensionality, leading to more efficient $\beta$-VAE training. Then, GPR maps the physical ROM input variables to the $\beta$-VAE latent space points. Finally, this work includes a final fine-tuning process that reduces the dependence on the $\beta$ hyperparameter, allowing easier model tuning. This integration and the fine-tuning process improve the accuracy and generalization capability of the model, marking a significant advancement in aerodynamic surrogate modeling using autoencoders.
@@ 2 Database and test case: XRF1 wing
The selected test case comprises a database of CFD simulations of the XRF1 model. This model, provided by Airbus™, is a research test case to showcase various technologies applied to long-range widebody aircrafts. The database was developed within the AD/AG60 research project of the Group for Aeronautical Research and Technology in Europe (GARTEUR). Further information on the XRF1 research model can be found in Pattinson and Herring [52].
@@
Aerodynamic data were derived from Reynolds Averaged Navier-Stokes (RANS) simulations of the complete aircraft model to ensure realistic conditions through the interaction of various aerodynamic subsystems. The dataset focuses on the dimensionless pressure coefficient ($C_{p}$) distribution over the wing for different flight conditions (see an example in Figure 1b). It comprises 435 distinct flight conditions computed using the DLR TAU solver [53] at a fixed Reynolds number ($Re=2.5\times 10^{7}$). The flight conditions span the entire flight envelope of the aircraft, with Mach numbers ($M$) ranging from 0.5 to 0.96 and angles of attack ($\alpha$) from $0^{\circ}$ to $11.5^{\circ}$ (see Figure 1a).
@@
Figure 1: (a) Flight envelope with the tested flight conditions split in training set ( ) and test set ( ). (b) Example of $C_{p}$ distribution over the wing for $\left(\alpha,M\right)=\left(9.25^{\circ},0.7\right)$. Three planes are highlighted at wingspan percentages $\eta=0.1,0.5,0.9$    ( ). Note that the wing deformation is the same for all the cases.
@@
The 435 flow solutions in the database are randomly split into two sets: a training set (70% of the database, 305 cases) used to train the surrogate model [( ) in Figure 1a], and a test set (30% of the database, 130 cases) to evaluate the model’s ability to predict $C_{p}$ within the entire design space [( ) in Figure 1a]. The cases in the test set are excluded from the training process. To visualize the performance of the models, three cases from the test set are chosen: low complexity $\left(\alpha_{1},M_{1}\right)=\left(2.0^{\circ},0.504\right)$, intermediate complexity $\left(\alpha_{2},M_{2}\right)=\left(5.75^{\circ},0.704\right)$, and high complexity $\left(\alpha_{3},M_{3}\right)=\left(9.75^{\circ},0.904\right)$, the latter featuring shockwaves on the upper wing surface. Although the performance of the model will be evaluated in the whole test set, these three cases will be used to visualize the error upon prediction with the proposed models.
@@
The fixed load factor $n_{z}$ ($=L/W$, where $L$ is the lift and $W$ is the weight) of 2.5 poses an additional challenge for the prediction model due to significant wing deformation under high loads, as can be seen in Figure 1b. This condition represents the limit load factor for transport category aircraft per airworthiness regulations. The wing undergoes structural deformation at such conditions but returns to its normal shape once the load factor returns to $n_{z}=1$ (typical cruise condition). Aerodynamically, the deflection of the wing geometry prevents shockwave formation on the upper surface, though significant $C_{p}$ changes and separation regions occur within the flight envelope. This condition is critical for surrogate model development due to its relevance to certification requirements.
@@
Based on the load factor condition, we only focus on the upper wing surface since the lower surface does not challenge the predictive models, as previously seen in Castellanos et al. [23]. Additionally, discarding the lower surface avoids the numerical artifacts in the model due to irregularities in the mesh around the engine pylon. Consequently, each $C_{p}$ sample is defined on an unstructured mesh of $q=49574$ grid points. Each $C_{p}$ distribution is an observation (point) in the high-dimensional space $\mathbb{R}^{q}$, where each dimension (feature) corresponds to a $C_{p}$ point. Let $\mathbf{X}\in{\mathbb{R}}^{q\times n}$ represent the data matrix, where $n$ indicates the number of available flight conditions (that is, 435). Each column vector, $\mathbf{x}_{i}\in\mathbb{R}^{q}$, corresponds to $C_{p,i}$ for $i=1,\ldots,n$. The dataset in $\mathbf{X}$ is inherently complex, challenging the extraction of a reduced set of meaningful coordinates that capture the main flow characteristics. Hereafter, data samples are denoted as $\mathbf{x}$, the dataset as $\mathbf{X}$, autoencoder reconstructions as $\mathbf{\tilde{x}}$, and surrogate model predictions as $\mathbf{\hat{x}}$. For clarity in the figures, the pressure distribution coefficient and its reconstructions and predictions are denoted as $C_{p}$, $\tilde{C}_{p}$, and $\hat{C}_{p}$, respectively.
@@ 3.1 $\beta-$variational autoencoder
Autoencoders are a recent field of research focused on using neural networks to derive meaningful, low-dimensional representations from high-dimensional data, leveraging the nonlinear nature of neural networks and their versatility in handling various types of data. They are an unsupervised learning strategy that aims to learn a compressed representation (latent space) of the training data by encoding it into a lower-dimensional space through an encoder network and, then reconstructing the input from this latent representation using a decoder network.
We denote as $\mathbf{z}_{i}$ the $i$-th component of $\mathbf{Z}\in\mathcal{Z}$, being $\mathcal{Z}\subset\mathbb{R}^{d}$ the latent space of dimensionality $d$, which is significantly smaller than the input dimension ($d\ll q$). The encoder network maps the input data, $\mathbf{x}$ in the data space $\mathcal{X}\subset\mathbb{R}^{q\times n}$, to the latent space ($E:\mathcal{X}\rightarrow\mathcal{Z}$) and the decoder network reconstructs the input data from the latent representation ($D:\mathcal{Z}\rightarrow\mathcal{X}$). The data is compressed into the latent space in the bottleneck between the encoder and decoder networks. Let $\tilde{\mathbf{x}}=(D\circ E)(\mathbf{x})$ be the output of the autoencoder, which is trained to minimize a reconstruction loss function $\mathcal{L}_{rec}$, to drive the reconstructed output data to match the input data. The standard autoencoder loss is given by the mean squared error (MSE),
$\mathcal{L}_{rec}=\frac{1}{q}\lVert\mathbf{x}-(D\circ E)(\mathbf{x})\rVert_{2}^{2},$ (1)
@@
where $\odot$ refers to the element-wise product, and $\mathcal{N}(0,1)$ refers to a Gaussian distribution with a mean of 0 and standard deviation of 1.
@@
In this paper, using $\beta$-VAEs to map pressure distribution fields enables the model to leverage a nonlinear dimensionality reduction method with increased fidelity over a classical linear method.